# Simulated Laser Cooling and Magneto-Optical Trapping of Group IV Atoms

### v1

@@ V White Light Slowing
The WLS parameters that we determined to provide effective slowing of Sn atoms for capture into the MOT are shown in Table 4. The spectral profile of slowing light is shown in Fig. 4, and a plot of the atomic deceleration is shown in the SM [45]. The short excited-state lifetime of the optical cycling transition in atomic Sn makes it possible to apply a high scattering force to the atoms. We find that a slowing distance of $z_{\text{MOT}}=500\,\text{mm}$ is sufficient for effective WLS. This is shorter than in typical molecular MOT setups [29, 30, 33], resulting in a larger solid angle subtended by the MOT region and hence greater slowing efficiency [74].
Table 4: Slowing laser parameters used for simulations of white light slowing of Sn atoms. We assume a converging slowing beam with $w=7.5$ mm ($1/e^{2}$ intensity radius) at the MOT center ($z=500$ mm), half-angle $\theta^{\prime}=4.5$ mrad, and linear polarization. $P$ indicates the slowing laser power, $\Delta_{s}$ the slowing laser carrier frequency detuning, $\Omega_{i}$ the modulation frequency of each EOM, and $\beta_{i}$ the modulation depth of each EOM.
P (mW) $\Delta_{s}$ ($\Gamma$) $\Omega_{1}$ ($\Gamma$) $\Omega_{2}$ ($\Gamma$) $\beta_{1}$ (rad) $\beta_{2}$ (rad)
300 11.75-11.75 1.5 3.4 1.5 1.5
@@ VI.1 Compressed MOT
Figure 6: Simulated compression sequence for the Sn red MOT. Arrows indicate the progression of the sequence. The temperature $T$ and cloud size $\sigma$ are reduced by factors of 30 and 4, respectively, as laser power $P$ is reduced, laser detuning is decreased, and B-field gradient is increased. We anticipate a compression sequence time of $\sim\!25$ ms, commensurate with what is typically used in MOT experiments [86].
@@ VI.2 Conveyor Belt Blue-Detuned MOT
Table 5: Parameters used for Sn CB MOT simulations (See Fig. 7). Laser power $P_{i}$ and detuning $\Delta_{i}$ are given for each polarization component $\hat{p}_{i}$ in the dual frequency scheme. $\partial B/\partial z$ is the axial B-field gradient. The Gaussian laser beam size ($1/e^{2}$ intensity radius) is $w_{\text{MOT}}=7$ $m$ in both stages. $t_{D}$ is the compression time; $T$ and $\sigma$ are the temperature and cloud size.
Label $\partial B/\partial z$ (G/cm) Laser Parameters Results
$P_{i}$ (mW) $\Delta_{i}(\Gamma)$ $\hat{p}_{i}$ $t_{D}$ (ms) $T$ ($\mu\text{K}$) $\sigma$ ($\mu\text{m}$)
Stage 1 30 172 +3.2+3.2 $\sigma^{-}$ 10 75 20
172 +3.0+3.0 $\sigma^{+}$
Stage 2 40 114 +4.2+4.2 $\sigma^{-}$ 15 15 10
114 +4.0+4.0 $\sigma^{+}$
@@ VII.3 Summary
Acknowledgements
We thank Silviu M. Udrescu for discussions about the applications of laser-cooled tin atoms for precision measurement. G.Z. and Q.W. acknowledge support from the Air Force Office of Scientific Research. M.V. acknowledges support from the Natural Sciences and Engineering Research Council of Canada.
Supplemental Material: Simulated Laser Cooling and Magneto-Optical Trapping of Group IV Atoms
SM I Calculating the Metastable $s^{2}p^{2}\,{{}^{3}}P_{1}$ State Lifetime
@@ Acknowledgements
$\hat{\mu}=-\frac{\mu_{B}}{\hbar}(g_{L}\mathbf{L}+g_{S}\mathbf{S}).$ (S1)
@@
$\Gamma_{\rm M1}=\frac{\omega_{0}^{3}}{3\pi\varepsilon_{0}\hbar c^{5}}\frac{|\langle g\|\hat{\mu}\|e\rangle|^{2}}{2J^{\prime}+1}.$ (S2)
@@
$\langle L,S,J\|\mathbf{L}\|L^{\prime},S^{\prime},J^{\prime}\rangle=\delta_{S,S^{\prime}}(-1)^{L^{\prime}+S+J^{\prime}+1}\sqrt{(2J+1)(2J^{\prime}+1)}\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}\langle L\|\mathbf{L}\|L^{\prime}\rangle,$ (S3)
@@
$\langle L,S,J\|\mathbf{S}\|L^{\prime},S^{\prime},J^{\prime}\rangle=\delta_{L,L^{\prime}}(-1)^{L+S^{\prime}+J+1}\sqrt{(2J+1)(2J^{\prime}+1)}\begin{Bmatrix}S^{\prime}&J^{\prime}&L\\ J&S&1\end{Bmatrix}\langle S\|\mathbf{S}\|S^{\prime}\rangle.$ (S4)
@@
$\Gamma_{\mathrm{M1}}$ $=\frac{\mu_{B}^{2}\omega_{0}^{3}}{3\pi\varepsilon_{0}\hbar c^{5}}\frac{1}{2J^{\prime}+1}\bigg{(}\delta_{S,S^{\prime}}\delta_{L,L^{\prime}}(-1)^{L^{\prime}+S+J^{\prime}+1}\sqrt{(2J+1)(2J^{\prime}+1)}$
$\quad\times\sqrt{L(L+1)(2L+1)}\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}+2\delta_{L,L^{\prime}}\delta_{S,S^{\prime}}(-1)^{L+S^{\prime}+J+1}$
$\quad\times\sqrt{(2J+1)(2J^{\prime}+1)}\sqrt{S(S+1)(2S+1)}\begin{Bmatrix}S^{\prime}&J^{\prime}&L\\ J&S&1\end{Bmatrix}\bigg{)}^{2}.$ (S5)
@@
Table S1: Parameters of interest for the M1 transition $s^{2}p^{2}\,{{}^{3}}P_{1}\leftrightarrow s^{2}p^{2}\,{{}^{3}}P_{0}$ for all Group IV elements. $k$ denotes transition wavenumber, $\omega_{0}$ denotes transition energy, $\Gamma_{\rm M1}$ denotes natural linewidth, and $\tau_{g}$ denotes lifetime of the metastable state.
@@
SM II Further Details on Atomic Tin MOT Simulations
@@
$a(r)=\frac{1}{2v_{\text{max}}}\int_{-v_{\text{max}}}^{v_{\text{max}}}a(r,v)\,dv,$ (S6)
@@
$a(v)=\frac{1}{2r_{\text{max}}}\int_{-r_{\text{max}}}^{r_{\text{max}}}a(r,v)\,dr.$ (S7)
Here, $v_{\text{max}}$ and $r_{\text{max}}$ are the maximum velocity and displacement of atoms in the capture MOT, which we find to be $v_{\text{max}}\approx 8$ m/s and $r_{\text{max}}\approx 3$ mm. Plots of these cuts are shown in Fig. S1(a) and (b), indicating the trapping and cooling behavior of the capture MOT stage. The sign reversal in the $a(v)$ curve at low velocities is a hallmark of sub-Doppler heating, a signature behavior of Type-II MOTs.
Fig. S1(c) shows how the capture velocity $v_{c}$ is determined, as described in the main text. We find that for velocities up to $v_{c}\sim 28.5$ m/s, atoms entering the MOT are slow enough to be captured. Captured atoms undergo a periodic, circular-like trajectory in phase-space, centered about the origin. This is indicative of sub-Doppler heating while trapped, due to the Type-II optical cycling transition.
Fig. S1(d) shows an example phase-space plot illustrative of how the temperature and size of the simulated capture MOT are determined. An atom is initialized at the phase-space origin and its trajectory propagated using the calculated acceleration $a(r,v)$. To account for random photon kicks from spontaneous emission during optical cycling, we give the atom a momentum kick in a random direction during each scattering event, which occurs with a rate $\sim P_{\rm exc}\Gamma$. The in-MOT trajectory with photon scattering is propagated for 50 ms to allow all observables to reach equilibrium. The temperature and size for each trajectory are calculated by computing the rms velocity and displacement in the last 10 ms of evolution. We average these values over 10 independent, stochastic trajectories to obtain the reported values for temperature $T$ and cloud size $\sigma$.
Figure S1: Analysis of the simulated Sn capture MOT. (a) Acceleration $a$ vs displacement $r$. The curve shows positional restoring behavior to the center of the trap. (b) Acceleration $a$ vs velocity $v$. The curve shows strong Doppler cooling leading to velocity damping for high velocities, as well as strong sub-Doppler heating at low velocities. (c) Phase space trajectory of Sn atoms entering the MOT region from the source beam. The arrowheads indicate the direction of time. Capture is possible for velocities up to $v_{c}\sim\!28.5$ m/s. (d) In-MOT trajectory of Sn atoms in the capture MOT, accounting for spontaneous photon emission due to optical cycling. The arrowhead indicates the direction of time. By computing the rms velocity and displacement averaged over the last 10 ms of the trajectory, we find the MOT temperature $T\sim\!225$ mK and cloud size $\sigma\sim\!1.9$ $m$.
@@
SM III Further Details on Tin White Light Slowing Simulations
@@
$e^{i\omega t+i\beta_{1}\sin(\Omega_{1}t)+i\beta_{2}\sin(\Omega_{2}t)}=e^{i\omega t}\left(J_{0}(\beta_{1})+\sum_{k=1}^{\infty}J_{k}(\beta_{1})e^{ik\Omega_{1}t}+\sum_{k=1}^{\infty}(-1)^{k}J_{k}(\beta_{1})e^{-ik\Omega_{1}t}\right)\times\\ \left(J_{0}(\beta_{2})+\sum_{k=1}^{\infty}J_{k}(\beta_{2})e^{ik\Omega_{2}t}+\sum_{k=1}^{\infty}(-1)^{k}J_{k}(\beta_{2})e^{-ik\Omega_{2}t}\right),$ (S8)
@@
After applying this broadened laser spectrum in the WLS simulations, we solve the OBEs and obtain the longitudinal deceleration $a_{z}(v_{z},r)$. We incorporate a converging slowing beam with half-angle $\theta^{\prime}=4.5$ mrad by solving the OBEs for $a_{z}(v_{z},r)$ as a function of longitudinally-varying laser intensity $s(z)$. Additionally, we take the transverse deceleration due to the converging slowing beam to be $a_{\perp}=a_{z}\tan\theta^{\prime}$. Interpolation across $z$ values from $z=0$ mm to $z=500$ mm enables us to obtain deceleration curves $a_{z}(v_{z},r,z)$, which are shown in Fig. S2. The calculated $a_{z}(v_{z},r,z)$ and $a_{\perp}(v_{z},r,z)$ are used to solve the differential equations for the trajectory of Sn atoms as they are slowed from source to MOT region.
Figure S2: Three-dimensional plot of the longitudinal deceleration $a_{z}(v_{z},z)$ imparted on Sn atoms due to white light slowing. Profiles for three different transverse displacements from the atomic beam axis $\hat{z}$ are shown: $r=0$ mm (blue), $r=3$ mm (green), and $r=6$ mm (brown). These show the decrease in slowing force as $r$ increases. The longitudinal intensity gradient due to the converging slowing beam is apparent as the peak value of $a_{z}$ gets increasingly negative for decreasing $z$.

### v2

@@ V White Light Slowing
The WLS parameters that we determined to provide effective slowing of Sn atoms for capture into the MOT are shown in Table 4. The spectral profile of slowing light is shown in Fig. 4, and a plot of the atomic deceleration is shown in Appendix C. The short excited-state lifetime of the optical cycling transition in atomic Sn makes it possible to apply a high scattering force to the atoms. We find that a slowing distance of $z_{\text{MOT}}=500\,\text{mm}$ is sufficient for effective WLS. This is shorter than in typical molecular MOT setups [96, 22, 105], resulting in a larger solid angle subtended by the MOT region and hence greater slowing efficiency [55].
Table 4: Slowing laser parameters used for simulations of white light slowing of Sn atoms. We assume a converging slowing beam with $w=7.5$ mm ($1/e^{2}$ intensity radius) at the MOT center ($z=500$ mm), half-angle $\theta^{\prime}=4.5$ mrad, and linear polarization. $P$ indicates the slowing laser power, $\bar{s}$ the saturation parameter spatially averaged along the slowing beamline axis, $\Delta_{s}$ the slowing laser carrier frequency detuning, $\Omega_{i}$ the modulation frequency of each EOM, and $\beta_{i}$ the modulation depth of each EOM.
$P$ (mW) $\bar{s}$ $\Delta_{s}$ ($\Gamma$) $\Omega_{1}$ ($\Gamma$) $\Omega_{2}$ ($\Gamma$) $\beta_{1}$ (rad) $\beta_{2}$ (rad)
300 3.3 11.75-11.75 1.5 3.4 1.5 1.5
@@ VI.1 Compressed MOT
Figure 6: Simulated compression sequence for the Sn red MOT. Arrows indicate the progression of the sequence. The temperature $T$ and cloud size $\sigma$ are reduced by factors of 30 and 4, respectively, as laser power $P$ is reduced ($P_{\text{capture}}$ is the laser power used for capture MOT stage), laser detuning is decreased, and B-field gradient is increased. We anticipate a compression sequence time of $\sim\!25$ ms, commensurate with what is typically used in MOT experiments [48].
@@ VI.2 Conveyor Belt Blue-Detuned MOT
Table 5: Parameters used for Sn CB MOT simulations (See Fig. 7). Laser power $P_{i}$, saturation parameter $s_{i}$, and detuning $\Delta_{i}$ are given for each polarization component $\hat{p}_{i}$ (for the given polarization, we assume same beam wave vector $\hat{k}_{i}$ direction as in previous MOT stages, see Table 3) in the dual frequency scheme. $\partial B/\partial z$ is the axial B-field gradient. The Gaussian laser beam size ($1/e^{2}$ intensity radius) is $w_{\text{MOT}}=7$ $m$ in both stages. $t_{D}$ is the compression time; $T$ and $\sigma$ are the temperature and cloud size.
Label $\frac{\partial B}{\partial z}$ (G/cm) Laser Parameters Results
$P_{i}$ (mW) $s_{i}$ $\Delta_{i}(\Gamma)$ $\hat{p}_{i}$ $t_{D}$ (ms) $T$ ($\mu\text{K}$) $\sigma$ ($\mu\text{m}$)
Stage 1 30 172 1.5 +3.2+3.2 $\sigma^{-}$ 10 75 20
172 1.5 +3.0+3.0 $\sigma^{+}$
Stage 2 40 114 1.0 +4.2+4.2 $\sigma^{-}$ 15 15 10
114 1.0 +4.0+4.0 $\sigma^{+}$
@@ VIII Summary
Acknowledgements.
We thank Silviu M. Udrescu for discussions about the applications of laser-cooled tin atoms for precision measurement. G.Z. and Q.W. acknowledge support from the Air Force Office of Scientific Research (AFOSR MURI Grant No. FA9550-21-1-0069). M.V. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC).
Data Availability
The data that support the findings of this article are openly available [107].
Appendix A Calculating the Metastable $s^{2}p^{2}\,{{}^{3}}\!P_{1}$ State Lifetime
@@
$\hat{\mu}=-\frac{\mu_{B}}{\hbar}(g_{L}\mathbf{L}+g_{S}\mathbf{S}).$ (2)
@@
$\Gamma_{\rm M1}=\frac{\omega_{0}^{3}}{3\pi\varepsilon_{0}\hbar c^{5}\!}\frac{|\langle g\|\hat{\mu}\|e\rangle|^{2}}{2J^{\prime}+1}.$ (3)
@@
$\langle L,S,J\|\mathbf{L}\|L^{\prime},S^{\prime},J^{\prime}\rangle=\delta_{S,S^{\prime}}(-1)^{L^{\prime}+S+J^{\prime}+1}\sqrt{(2J+1)(2J^{\prime}+1)}\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}\langle L\|\mathbf{L}\|L^{\prime}\rangle,$ (4)
@@
$\langle L,S,J\|\mathbf{S}\|L^{\prime},S^{\prime},J^{\prime}\rangle=\delta_{L,L^{\prime}}(-1)^{L+S^{\prime}+J+1}\sqrt{(2J+1)(2J^{\prime}+1)}\begin{Bmatrix}S^{\prime}&J^{\prime}&L\\ J&S&1\end{Bmatrix}\langle S\|\mathbf{S}\|S^{\prime}\rangle.$ (5)
@@
$\Gamma_{\mathrm{M1}}=\frac{\mu_{B}^{2}\omega_{0}^{3}}{3\pi\varepsilon_{0}\hbar c^{5}\!}\,\frac{|\mathcal{M}|^{2}}{2J^{\prime}+1},$ (6)
where
$\mathcal{M}$ $=\delta_{S,S^{\prime}}\delta_{L,L^{\prime}}(-1)^{L^{\prime}+S+J^{\prime}+1}\sqrt{(2J+1)(2J^{\prime}+1)}\sqrt{L(L+1)(2L+1)}\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}$ (7)
$\quad+2\,\delta_{L,L^{\prime}}\delta_{S,S^{\prime}}(-1)^{L+S^{\prime}+J+1}\sqrt{(2J+1)(2J^{\prime}+1)}\sqrt{S(S+1)(2S+1)}\begin{Bmatrix}S^{\prime}&J^{\prime}&L\\ J&S&1\end{Bmatrix}.$
@@
Table 6: Parameters of interest for the M1 transition $s^{2}p^{2}\,{{}^{3}}\!P_{1}\leftrightarrow s^{2}p^{2}\,{{}^{3}}\!P_{0}$ for all Group IV elements. $k$ denotes transition wavenumber, $\omega_{0}$ denotes transition energy, $\Gamma_{\rm M1}$ denotes natural linewidth, and $\tau_{g}$ denotes lifetime of the metastable state.
@@
Appendix B Further Details on Atomic Tin MOT Simulations
@@
$a(r)=\frac{1}{2v_{\text{max}}}\int_{-v_{\text{max}}}^{v_{\text{max}}}a(r,v)\,dv,$ (8)
@@
$a(v)=\frac{1}{2r_{\text{max}}}\int_{-r_{\text{max}}}^{r_{\text{max}}}a(r,v)\,dr.$ (9)
Here, $v_{\text{max}}$ and $r_{\text{max}}$ are the maximum velocity and displacement of atoms in the capture MOT, which we find to be $v_{\text{max}}\approx 8$ m/s and $r_{\text{max}}\approx 3$ mm. Plots of these cuts are shown in Fig. 8(a) and (b), indicating the trapping and cooling behavior of the capture MOT stage. The sign reversal in the $a(v)$ curve at low velocities is a hallmark of sub-Doppler heating, a signature behavior of Type-II MOTs.
Fig. 8(c) shows how the capture velocity $v_{c}$ is determined, as described in the main text. We find that for velocities up to $v_{c}\sim 28.5$ m/s, atoms entering the MOT are slow enough to be captured. Captured atoms undergo a periodic, circular-like trajectory in phase-space, centered about the origin. This is indicative of sub-Doppler heating while trapped, due to the Type-II optical cycling transition.
Fig. 8(d) shows an example phase-space plot illustrative of how the temperature and size of the simulated capture MOT are determined. An atom is initialized at the phase-space origin and its trajectory propagated using the calculated acceleration $a(r,v)$. To account for random photon kicks from spontaneous emission during optical cycling, we give the atom a momentum kick in a random direction during each scattering event, which occurs with a rate $\sim P_{\rm exc}\Gamma$. The in-MOT trajectory with photon scattering is propagated for 50 ms to allow all observables to reach equilibrium. The temperature and size for each trajectory are calculated by computing the rms velocity and displacement in the last 10 ms of evolution. We average these values over 10 independent, stochastic trajectories to obtain the reported values for temperature $T$ and cloud size $\sigma$.
Figure 8: Analysis of the simulated Sn capture MOT. (a) Acceleration $a$ vs displacement $r$. The curve shows positional restoring behavior to the center of the trap. (b) Acceleration $a$ vs velocity $v$. The curve shows strong Doppler cooling leading to velocity damping for high velocities, as well as strong sub-Doppler heating at low velocities. (c) Phase space trajectory of Sn atoms entering the MOT region from the source beam. The arrowheads indicate the direction of time. Capture is possible for velocities up to $v_{c}\sim\!28.5$ m/s. (d) In-MOT trajectory of Sn atoms in the capture MOT, accounting for spontaneous photon emission due to optical cycling. The arrowhead indicates the direction of time. By computing the rms velocity and displacement averaged over the last 10 ms of the trajectory, we find the MOT temperature $T\sim\!225$ mK and cloud size $\sigma\sim\!1.9$ $m$.
@@
Appendix C Further Details on Tin White Light Slowing Simulations
@@
$e^{i\omega t+i\beta_{1}\sin(\Omega_{1}t)+i\beta_{2}\sin(\Omega_{2}t)}=e^{i\omega t}\left(J_{0}(\beta_{1})+\sum_{k=1}^{\infty}J_{k}(\beta_{1})e^{ik\Omega_{1}t}+\sum_{k=1}^{\infty}(-1)^{k}J_{k}(\beta_{1})e^{-ik\Omega_{1}t}\right)\times\\ \left(J_{0}(\beta_{2})+\sum_{k=1}^{\infty}J_{k}(\beta_{2})e^{ik\Omega_{2}t}+\sum_{k=1}^{\infty}(-1)^{k}J_{k}(\beta_{2})e^{-ik\Omega_{2}t}\right),$ (10)
@@
After applying this broadened laser spectrum in the WLS simulations, we solve the OBEs and obtain the longitudinal deceleration $a_{z}(v_{z},r)$. We incorporate a converging slowing beam with half-angle $\theta^{\prime}=4.5$ mrad by solving the OBEs for $a_{z}(v_{z},r)$ as a function of longitudinally-varying laser intensity $s(z)$. Additionally, we take the transverse deceleration due to the converging slowing beam to be $a_{\perp}=a_{z}\tan\theta^{\prime}$. Interpolation across $z$ values from $z=0$ mm to $z=500$ mm enables us to obtain deceleration curves $a_{z}(v_{z},r,z)$, which are shown in Fig. 9. The calculated $a_{z}(v_{z},r,z)$ and $a_{\perp}(v_{z},r,z)$ are used to solve the differential equations for the trajectory of Sn atoms as they are slowed from source to MOT region.
Figure 9: Three-dimensional plot of the longitudinal deceleration $a_{z}(v_{z},z)$ imparted on Sn atoms due to white light slowing. Profiles for three different transverse displacements from the atomic beam axis $\hat{z}$ are shown: $r=0$ mm (blue), $r=3$ mm (green), and $r=6$ mm (brown). These show the decrease in slowing force as $r$ increases. The longitudinal intensity gradient due to the converging slowing beam is apparent as the peak value of $a_{z}$ gets increasingly negative for decreasing $z$.